# eprint Nisho-1-2025

### v2

@@ Abstract
We propose a new method for detecting dark matter axions using a resonant cavity coupled with a quantum Hall system. When a sample exhibiting quantum Hall effect is placed inside the cavity and the cavity is tuned to resonance, two-dimensional electrons absorb the amplified radiation, leading to a rise in the sample’s temperature. By monitoring this temperature increase, the mass $m_{a}$ of the axion can be inferred. As an example, consider a GaAs sample with small thickness $d=0.1\,\mu\mathrm{m}$ and its heat capacity $C_{s}$ at temperature $T=20\,\mathrm{mK}$. According to the QCD axion model, because the energy flux of the incoming radiation is $P_{ra}\sim 10^{-18}g_{\gamma}^{2}\mbox{W cm}^{-2}(10^{-5}\mbox{eV}/m_{a})(B/1.5\times 10^{5}\mbox{Gauss})^{2}(\rho_{d}/0.3\rm GeVcm^{-3})$ at the resonance with $g_{\gamma}(\rm KSVZ)=-0.96$ and $g_{\gamma}(\rm DFSZ)=0.37$, the temperature increase $(P_{ra}/C_{s})\times\tau$ is approximately $\simeq 0.25(\mbox{mK})(\tau/10\mbox{ms})\,g_{\gamma}^{2}\,(20\mbox{mK}/T)^{3}(10^{-5}\mbox{eV}/m_{a})(B/1.5\times 10^{5}\mbox{Gauss})^{2}(0.1\mu\mbox{m}/d)$, where $\tau$ is the time constant associated with the heat dissipation into thermal bath.
@@
The temperature increase due to energy absorption depends on the heat capacity of the semiconductor sample. The total heat capacity $C$ is the sum of the semiconductor heat capacity $C_{s}$ and the 2D electron contribution $C_{e}$. By appropriately tuning the Fermi energy, the contribution from the 2D electrons $C_{e}$ can be made negligible compared to $_{s}; CC$$\simeq C_{s}$. At low temperatures (e.g., 1010mK), the semiconductor heat capacity $C_{s}$ is dominated by phonons and follows the relation $C_{s}\propto T^{3}$, making the temperature increase highly sensitive to the temperature $T$. The lower the temperature, the greater the temperature rise for a given amount of absorbed energy. In our analysis, we assume a temperature of T=20T=20mK, which is relatively easy to achieve experimentally.
@@
$\frac{P_{ra}}{C}\simeq\frac{0.025\,\mathrm{mK}}{\mathrm{ms}}\times g_{\gamma}^{2}\left(\frac{20\,\mathrm{mK}}{T}\right)^{3}\Big(\frac{10^{-5}\mbox{eV}}{m_{a}}\Big)\left(\frac{B_{t}}{1.5\times 10^{5}\,\mathrm{Gauss}}\right)^{2}\left(\frac{0.1\,\mu\mathrm{m}}{d}\right)\Big(\frac{\rho_{d}}{0.3\rm GeVcm^{-3}}\Big),$ (1)
@@
It is important to note that this temperature rise rate is independent of the surface area $S$ of the sample, since both the absorbed power $P_{ra}\propto S$ and the heat capacity $C_{s}\propto S$. The form of semiconductor sample is rectangular with surface area $S$ and thickness $d$.
Finally, because the absorbed energy dissipates into the thermal bath, the actual observed temperature increase $\Delta T$ depends on the time constant $\tau$ of the dissipation. For example, with the time constant $\tau=10$ms, we have
$\Delta T\simeq 0.25\,\mathrm{mK}\times\left(\frac{\tau}{10\,\mathrm{ms}}\right)g_{\gamma}^{2}\left(\frac{20\,\mathrm{mK}}{T}\right)^{3}\Big(\frac{10^{-5}\mbox{eV}}{m_{a}}\Big)\Big(\frac{B_{t}}{1.5\times 10^{5}\mbox{Gauss}}\Big)^{2}\left(\frac{0.1\,\mu\mathrm{m}}{d}\right)\Big(\frac{\rho_{d}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
@@
$\Delta T\simeq 0.5\text{mK}\left(\frac{\tau}{10\,\text{ms}}\right)\,\Big(\frac{g_{a\gamma\gamma}}{10^{-14}\text{GeV}^{-1}}\Big)^{2}\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)^{3}\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$ (3)
It is favorable to cover the semiconductor sample with a transparent thin sheet possessing much small heat capacity and much small heat conductance for $\tau$ to be larger than 1010ms.
We begin by reviewing the fundamental properties of the quantum Hall state in 2D electron systems. We then demonstrate how microwave radiation, generated by axions in the presence of an external magnetic field $B_{t}$, is absorbed by the 2D electrons. Our analysis shows that the entire flux of axion-induced microwave radiation is fully absorbed in the quantum Hall state. Through out the paper, the temperature is of the order of 1010mK so that chemical potential is nearly equal to the Fermi energy $E_{f}$. We use physical units, Boltzmann constant $k_{B}=1$, light velocity $c=1$ and $\hbar=1$.
@@
In order to discuss the absorption of the axion induced radiation in quantum Hall states, we explain electromagnetic coupling with axionaxion1 ; axion2 ; axion3 and smallness of the coupling. The coupling $g_{a\gamma\gamma}$ between axion $a(t,\vec{x})$ and electromagnetic field is given by $L_{a\gamma\gamma}=g_{a\gamma\gamma}a(t,\vec{x})\vec{E}\cdot\vec{B}$ with electric $\vec{E}$ and magnetic $\vec{B}$ fields. The coupling $g_{a\gamma\gamma}$ is rewritten such that $g_{a\gamma\gamma}=g_{\gamma}\alpha/f_{a}\pi$, where fine structure constant $\alpha\simeq 1/137$ and axion decay constant $f_{a}$ satisfying the relation $m_{a}f_{a}\simeq 6\times 10^{-6}\rm eV\times 10^{12}$GeV in the QCD axion. The parameter $g_{\gamma}$ depends on the axion model, i.e. $g_{\gamma}(\rm DFSZ)\simeq 0.37$ for DFSZ modeldfsz ; dfsz1 and $g_{\gamma}(\rm KSVZ)\simeq-0.96$ for KSVZ modelksvz ; ksvz1 . Hereafter we only consider QCD axion model.
@@
To detect the amplified radiation, we utilize a quantum Hall system placed inside a resonant cavity. When a semiconductor sample exhibiting the quantum Hall effect absorbs axion induced microwaves amplified by resonance, its temperature increases. We can determine the presence of the axion and its mass with the detection of the temperature increase. This provides a new method for axion detection.
@@
To make our discussion more concrete, we consider a resonant cavity with parallel slabs, as shown in Fig.(2), following our previous work iwazaki01 . Let the distance between the two parallel, facing slabs be $l$, and assume that an external magnetic field $\vec{B}_{t}=(0,0,B_{t})$ is applied parallel to the slabs. By adjusting the distance $l$, the resonance condition can be satisfied.
@@
$\frac{\Delta N\tau_{t}^{-1}}{3.8Sm_{a}A_{0}^{\prime 2}}\simeq 6.7\times 10^{5}\Big(\frac{S}{1\rm cm^{2}}\Big)\Big(\frac{B}{10^{5}\rm Gauss}\Big)\Big(\frac{m_{a}}{10^{-5}\rm eV}\Big)^{2}\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)^{2}A^{2}\exp\Big(-2\Big(\frac{E_{f}-E_{c}}{\Delta E}\Big)^{2}\Big)>1.$ (14)
@@
Namely, all photons $3.8Sm_{a}A_{0}^{\prime 2}$ incoming are absorbed within a time $\tau_{t}$, because there are sufficiently number of electrons $\Delta N$ which are able to absorb the photons within the time $\tau_{t}$. The condition in eq(14) holds even for external microwave irradiation with frequency $m_{a}$ and any power. Then, it apparently seems that incoming microwaves with any power are fully absorbed.
@@
$\tau_{t}\sim 10^{-3}\text{ s}\left(\frac{1\text{ cm}^{2}}{S}\right)\left(\frac{m_{a}}{10^{-5}\text{ eV}}\right)\left(\frac{10^{5}\text{ Gauss}}{B}\right)^{2}\left(\frac{1}{A^{2}}\right)\left(\frac{0.3\,\text{GeV}/\text{cm}^{3}}{\rho_{d}}\right)\exp\left(\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right),$ (15)
@@
$P_{ra}\sim 0.46SE_{a}^{2}\sim 10^{-30}\text{ W}\left(\frac{B}{10^{5}\text{ Gauss}}\right)^{2}\left(\frac{S}{1\text{cm}^{2}}\right)\left(\frac{\rho_{d}}{0.3\,\text{GeV}/\text{cm}^{3}}\right),$ (16)
but it can be amplified to $\sim 10^{-18}\text{ W}$ using a properly designed resonant cavity.
@@
$P_{ra}\sim 0.46\left(\frac{E_{a}}{m_{a}\delta_{e}}\right)^{2}S\sim 0.9\times 10^{-18}\,\text{W}\,g_{\gamma}^{2}\frac{S}{\text{cm}^{2}}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV/cm}^{3}}\right),$ (20)
@@
$C_{s}\simeq 0.88\times 10^{7}\left(\frac{T}{20\,\text{mK}}\right)^{3}\left(\frac{S}{1\,\text{cm}^{2}}\right)\left(\frac{d}{0.1\,\mu\text{m}}\right).$ (22)
@@
$C_{e}\simeq 3.4\times 10^{8}\left(\frac{S}{1\,\text{cm}^{2}}\right)\left(\frac{T}{20\,\text{mK}}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{1/2}\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right).$ (25)
Thus, when $\exp\left(-(E_{f}-E_{c})^{2}/\Delta E^{2}\right)<10^{-2}$, the dominant contribution to the total heat capacity is from the semiconductor, i.e., $C_{s}+C_{e}\approx C_{s}$.
@@
$10^{-3}\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)^{-1}A^{-1}<\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right)<10^{-2}$ (26)
@@
$\dot{T}=\frac{P_{ra}}{C_{s}(T)}\simeq\frac{7.2\,\text{mK}}{\text{s}}\,g_{\gamma}^{2}\left(\frac{20\,\text{mK}}{T}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}/\text{cm}^{3}}\right)$ (27)
where we take the thickness $d=0.1\,\mu\text{m}$ and temperature $T=20\,\text{mK}$, with an external magnetic field $B_{t}=1.5\times 10^{5}\,\text{Gauss}$, applied parallel to the sample layers. ( Even when smaller magnetic field $B_{t}=6\times 10^{4}$Gauss, $\dot{T}\simeq 0.032$mK/ms with lower temperature T=10T=10mK. )
@@
Thus, the final temperature increase due to axion-induced microwave absorption is approximately
$\Delta T(t=\infty)\simeq$ $0.72$ $\,\text{mK}\left(\frac{\tau}{100\,\text{ms}}\right)\,g_{\gamma}^{2}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)$ (30)
$\times$ $\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$
@@
$\Delta T(t=\infty)\simeq$ $5$ $\text{mK}\left(\frac{\tau}{100\,\text{ms}}\right)\,\Big(\frac{g_{a\gamma\gamma}}{10^{-14}\text{GeV}^{-1}}\Big)^{2}\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)$ (31)
$\times$ $\left(\frac{0.1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$
We show in Fig.(3) the excluded region of $g_{a\gamma\gamma}$ and $m_{a}$ indicated by dashed line when we do not observe $\Delta T$ equal to or larger than $0.5\text{mK}(\tau/10\text{ms})$ in the case of time constant $\tau=10$ms. If the time constant $\tau$ is less than 1010ms, for instance, 11ms, the excluded region is above the same dashed line indicated with $\Delta T=0.05\text{mK}(\tau/1\text{ms})$ as the line in the figure when we do not observe $\Delta T\geq 0.05$mK. In the figure, we have roughly depicted the excluded region by CAPPCAPP1 ; CAPP2 ; CAPP3 . We also show the lines associated with the models of QCD axion.
Figure 3: When we do not observe $\Delta T$ larger than $0.5\text{mK}(\tau/10\text{ms})$ with time constant $\tau=10$ms, the value of $g_{a\gamma\gamma}$ above the dashed line denoted with $0.5\text{mK}(\tau/10\text{ms})$ is excluded. We roughly show excluded region obtained by CAPP.
Therefore, we find that a large time constant $\tau>10\,\text{ms}$ is preferable to observe a measurable temperature increase $\Delta T$. To achieve such a large time constant, it is needed to cover the semiconductor sample with a transparent, thin sheet that has both lower thermal conductivity and lower heat capacity than those of the sample semiconductor. Additionally, minimizing the heat capacity $C_{s}$ by using a very thin sample (e.g., $d\leq 0.1\,\mu\text{m}$) is advantageous. Importantly, the result $\Delta T$ does not depend on the surface area $S$ of the sample since both $P_{ra}\propto S$ and $C_{s}\propto S$; $\Delta T\propto P_{ra}/C_{s}$.
@@
Our idea for the detection of dark matter axion is that we measure the increase of the temperature caused by the absorption of much weak microwaves generated by the axion. So it is favorable to use materials with much small heat capacity. In the paper we have proposed two dimensional electrons forming quantum Hall states as such a material. But the idea is also applicable for materials which are able to absorb the microwaves and have sufficiently small heat capacities for the temperature to increase. To find such a material is a subject in near future.
@@
Although the method for measuring the temperature rise is beyond the scope of this work, the predicted increase offers a direct and sensitive means for detecting the axion and determining its mass.

### v3

@@ Abstract
We propose a new method for detecting dark matter axions using a resonant cavity coupled with a quantum Hall system. When a sample exhibiting quantum Hall effect is placed inside the cavity and the cavity is tuned to resonance, two-dimensional electrons absorb the amplified radiation, leading to a rise in the sample’s temperature. By monitoring this temperature increase, the mass $m_{a}$ of the axion can be inferred. As an example, consider a GaAs sample with small thickness $d=1\,\mu\mathrm{m}$ and its heat capacity $C_{s}$ at temperature $T=20\,\mathrm{mK}$. According to the QCD axion model, because the energy flux of the incoming radiation is $P_{ra}\sim 10^{-18}g_{\gamma}^{2}\mbox{W cm}^{-2}(10^{-5}\mbox{eV}/m_{a})(B/1.5\times 10^{5}\mbox{Gauss})^{2}(\rho_{d}/0.3\rm GeVcm^{-3})$ at the resonance with $g_{\gamma}(\rm KSVZ)=-0.96$ and $g_{\gamma}(\rm DFSZ)=0.37$, the temperature increase $(P_{ra}/C_{s})\times t_{ob}$ is approximately $\simeq 0.72(\mbox{mK})(t_{ob}/1\text{s})\,g_{\gamma}^{2}\,(20\mbox{mK}/T)^{3}(10^{-5}\mbox{eV}/m_{a})(B/1.5\times 10^{5}\mbox{Gauss})^{2}(1\mu\text{m}/d)$, where $t_{ob}$ is the observation time. It must be smaller than a time constant $\tau$ associated with the heat dissipation into thermal bath. Such a large time constant should be realized by making thermal conductance much small.
@@
For larger axion masses $m_{a}$, the cavity volume scales as $m_{a}^{-3}$ and therefore becomes smaller, leading to a corresponding reduction in the axion-induced power, which also scales as $m_{a}^{-3}$. As a result, direct power measurements in conventional axion haloscopes become increasingly challenging at higher axion masses.
At resonance, the energy density of the amplified radiation scales as $m_{a}^{-1}$ and thus also decreases with increasing axion mass, but much more gradually than the axion-induced power. Consequently, the local energy absorbed by a small semiconductor sample is proportional to $m_{a}^{-1}$, and the resulting temperature rise rate follows the same scaling. This weaker dependence on $m_{a}$ makes temperature-based measurements a more effective approach for axion detection at higher masses. We note that the surface area of the sample is of the order of $(0.1\text{cm})^{2}$ or less. Only a fraction of induced radiations by the axion in the cavity is absorbed. Putting such a small sample in the resonant cavity, we can determine the presence of the axion and its mass with the detection of the temperature increase. This is an advantage over previous axion haloscopes.
@@
The temperature increase due to energy absorption depends on the heat capacity of the semiconductor sample. The total heat capacity $C$ is the sum of the semiconductor heat capacity $C_{s}$ and the 2D electron contribution $C_{e}$. By appropriately tuning the Fermi energy, the contribution from the 2D electrons $C_{e}$ can be made negligible compared to $_{s}; CC$$\simeq C_{s}$. Or simply by taking the volume, e.g. thickness of the semiconductor sample large, heat capacity $C_{s}$ can be larger than $C_{e}$. At low temperatures (e.g., 1010mK), the semiconductor heat capacity $C_{s}$ is dominated by phonons and follows the relation $C_{s}\propto T^{3}$, making the temperature increase highly sensitive to the temperature $T$. The lower the temperature, the greater the temperature rise for a given amount of absorbed energy. In our analysis, we assume a temperature of T=20T=20mK, which is relatively easy to achieve experimentally.
@@
$\frac{P_{ra}}{C}\simeq\frac{0.72\,\mathrm{mK}}{\mathrm{s}}\times g_{\gamma}^{2}\left(\frac{20\,\mathrm{mK}}{T}\right)^{3}\Big(\frac{10^{-5}\mbox{eV}}{m_{a}}\Big)\left(\frac{B_{t}}{1.5\times 10^{5}\,\mathrm{Gauss}}\right)^{2}\left(\frac{1\,\mu\mathrm{m}}{d}\right)\Big(\frac{\rho_{d}}{0.3\rm GeVcm^{-3}}\Big),$ (1)
@@
It is important to note that this temperature rise rate is independent of the surface area $S$ of the sample, since both the absorbed power $P_{ra}\propto S$ and the heat capacity $C_{s}\propto S$. The form of semiconductor sample is rectangular with surface area $S$ and thickness $d$. It apparently seems that the temperature is larger as the observation time $t_{ob}$ is longer, $\Delta T=P_{ra}/C\times t_{ob}$.
But, the absorbed energy dissipates into the thermal bath and the actual observed temperature increase $\Delta T$ depends on the time constant $\tau$ of the dissipation. However, the energy dissipation is very slow because the semiconductor sample is connected to thermal bath with small copper lines for measurement of electric current and small pedestal supposing the sample in dilution refrigerator. The surface area in contact with the thermal bath is much smaller than the one of the sample. Thus, the actual temperature increase is determined by the the observation time $t_{ob}$ as far as $t_{ob}<\tau$. As we discuss later, we may take $t_{ob}=1$ second and obtain for QCD axion,
$\Delta T\simeq 0.72\,\mathrm{mK}\times\left(\frac{t_{ob}}{1\,\mathrm{s}}\right)g_{\gamma}^{2}\left(\frac{20\,\mathrm{mK}}{T}\right)^{3}\Big(\frac{10^{-5}\mbox{eV}}{m_{a}}\Big)\Big(\frac{B_{t}}{1.5\times 10^{5}\mbox{Gauss}}\Big)^{2}\left(\frac{1\,\mu\mathrm{m}}{d}\right)\Big(\frac{\rho_{d}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
@@
$\Delta T\simeq 5\text{mK}\left(\frac{t_{ob}}{1\,\text{s}}\right)\,\Big(\frac{g_{a\gamma\gamma}}{10^{-14}\text{GeV}^{-1}}\Big)^{2}\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)^{3}\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{1\,\mu\text{m}}{d}\right)\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$ (3)
It would be not difficult to fabricate the sample with small copper lines and small pedestal for time constant $\tau$ of the energy dissipation to be sufficiently longer than the observation time, i.e. 11\,s.
We begin by reviewing the fundamental properties of the quantum Hall state in 2D electron systems. We then demonstrate how microwave radiation, generated by axions in the presence of an external magnetic field $B_{t}$, is absorbed by the 2D electrons. Our analysis shows that the entire radiation flux entering the sample is fully absorbed in the quantum Hall state. Through out the paper, the temperature is of the order of 1010mK so that chemical potential is nearly equal to the Fermi energy $E_{f}$. We use physical units, Boltzmann constant $k_{B}=1$, light velocity $c=1$ and $\hbar=1$.
@@
In order to discuss the absorption of the axion induced radiation in quantum Hall states, we explain electromagnetic coupling with axionaxion1 ; axion2 ; axion3 and smallness of the coupling. The coupling $g_{a\gamma\gamma}$ between axion $a(t,\vec{x})$ and electromagnetic field is given by $L_{a\gamma\gamma}=g_{a\gamma\gamma}a(t,\vec{x})\vec{E}\cdot\vec{B}$ with electric $\vec{E}$ and magnetic $\vec{B}$ fields. The coupling $g_{a\gamma\gamma}$ is rewritten such that $g_{a\gamma\gamma}=g_{\gamma}\alpha/f_{a}\pi$, where fine structure constant $\alpha\simeq 1/137$ and axion decay constant $f_{a}$ satisfying the relation $m_{a}f_{a}\simeq 6\times 10^{-6}\rm eV\times 10^{12}$GeV in the QCD axion. The parameter $g_{\gamma}$ depends on the axion model, i.e. $g_{\gamma}(\rm DFSZ)\simeq 0.37$ for DFSZ modeldfsz ; dfsz1 and $g_{\gamma}(\rm KSVZ)\simeq-0.96$ for KSVZ modelksvz ; ksvz1 . Hereafter we mainly consider QCD axion model, although our results are also presented with general $g_{a\gamma\gamma}$ independent of $m_{a}$.
@@
To detect the amplified radiation, we utilize a quantum Hall system placed inside a resonant cavity. When a semiconductor sample exhibiting the quantum Hall effect absorbs axion induced microwaves amplified by resonance, its temperature increases. The surface area of the sample is of the order of $(0.1\text{cm})^{2}$ or less so that only a fraction of induced radiations by the axion in the cavity is absorbed. Putting such a small sample in the resonant cavity, we can determine the presence of the axion and its mass with the detection of the temperature increase. This provides a new method for axion detection.
@@
To make our discussion more concrete, we consider a resonant cavity with parallel slabs, as shown in Fig.(2), following our previous work iwazaki01 . Let the distance between the two parallel, facing slabs be $l$, and assume that an external magnetic field $\vec{B}_{t}=(0,0,B_{t})$ is applied parallel to the slabs. By adjusting the distance $l$, the resonance condition can be satisfied. We put a semiconductor sample involving two dimensional electrons in quantum Hall state. The sample in the cavity is very small compared with the length $l$. The surface area is of the order of $(0.1\text{cm})^{2}$ or less.
@@
$\frac{\Delta N\tau_{t}^{-1}}{3.8Sm_{a}A_{0}^{\prime 2}}\simeq 6.7\times 10^{3}\Big(\frac{S}{\rm(0.1cm)^{2}}\Big)\Big(\frac{B}{10^{5}\rm Gauss}\Big)\Big(\frac{m_{a}}{10^{-5}\rm eV}\Big)^{2}\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)^{2}A^{2}\exp\Big(-2\Big(\frac{E_{f}-E_{c}}{\Delta E}\Big)^{2}\Big)>1.$ (14)
@@
Namely, all photons $3.8Sm_{a}A_{0}^{\prime 2}$ incoming to the semiconductor sample are absorbed within a time $\tau_{t}$, because there are sufficiently number of electrons $\Delta N$ which are able to absorb the photons within the time $\tau_{t}$. The condition in eq(14) holds even for external microwave irradiation with frequency $m_{a}$ and any power. Then, it apparently seems that incoming microwaves with any power are fully absorbed.
@@
$\tau_{t}\sim 10^{-1}\text{ s}\left(\frac{\text{(0.1cm)}^{2}}{S}\right)\left(\frac{m_{a}}{10^{-5}\text{ eV}}\right)\left(\frac{10^{5}\text{ Gauss}}{B}\right)^{2}\left(\frac{1}{A^{2}}\right)\left(\frac{0.3\,\text{GeV}/\text{cm}^{3}}{\rho_{d}}\right)\exp\left(\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right),$ (15)
@@
$P_{ra}\sim 0.46SE_{a}^{2}\sim 10^{-32}\text{ W}\left(\frac{B}{10^{5}\text{ Gauss}}\right)^{2}\left(\frac{S}{\text{(0.1cm)}^{2}}\right)\left(\frac{\rho_{d}}{0.3\,\text{GeV}/\text{cm}^{3}}\right),$ (16)
but it can be amplified to $\sim 10^{-20}\text{ W}$ using a properly designed resonant cavity.
@@
$P_{ra}\sim 0.46\left(\frac{E_{a}}{m_{a}\delta_{e}}\right)^{2}S\sim 0.9\times 10^{-20}\,\text{W}\,g_{\gamma}^{2}\Big(\frac{S}{\text{(0.1cm)}^{2}}\Big)\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV/cm}^{3}}\right),$ (20)
@@
$C_{s}\simeq 0.88\times 10^{6}\left(\frac{T}{20\,\text{mK}}\right)^{3}\left(\frac{S}{\text{(0.1cm)}^{2}}\right)\left(\frac{d}{1\,\mu\text{m}}\right).$ (22)
@@
$C_{e}\simeq 3.4\times 10^{6}\left(\frac{S}{\text{(0.1cm)}^{2}}\right)\left(\frac{T}{20\,\text{mK}}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{1/2}\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right).$ (25)
Thus, when $\exp\left(-(E_{f}-E_{c})^{2}/\Delta E^{2}\right)<10^{-1}$, the dominant contribution to the total heat capacity is from the semiconductor, i.e., $C_{s}+C_{e}\approx C_{s}$.
@@
$10^{-3}\Big(\frac{10^{-3}\rm eV}{\Delta E}\Big)^{-1}A^{-1}<\exp\left(-\left(\frac{E_{f}-E_{c}}{\Delta E}\right)^{2}\right)<10^{-1}$ (26)
@@
$\dot{T}=\frac{P_{ra}}{C_{s}(T)}\simeq\frac{0.72\,\text{mK}}{\text{s}}\,g_{\gamma}^{2}\left(\frac{20\,\text{mK}}{T}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}/\text{cm}^{3}}\right)$ (27)
where we take the thickness $d=1\,\mu\text{m}$ and temperature $T=20\,\text{mK}$, with an external magnetic field $B_{t}=1.5\times 10^{5}\,\text{Gauss}$, applied parallel to the sample layers. ( Even when smaller magnetic field $B_{t}=10^{5}$Gauss, $\dot{T}\simeq 0.32$mK/s with lower temperature T=20T=20mK. )
@@
[…]
Thus, the temperature rise we observe with the observation time $t_{ob}(<\tau)$ is approximately
$\Delta T(t=\infty)\simeq$ $0.72$ $\,\text{mK}\left(\frac{t_{ob}}{1\,\text{s}}\right)\,g_{\gamma}^{2}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)$ (30)
$\times$ $\left(\frac{1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$
@@
$\Delta T(t=\infty)\simeq$ $5$ $\text{mK}\left(\frac{t_{ob}}{1\,\text{s}}\right)\,\Big(\frac{g_{a\gamma\gamma}}{10^{-14}\text{GeV}^{-1}}\Big)^{2}\left(\frac{20\text{mK}}{T}\right)^{3}\left(\frac{10^{-5}\,\text{eV}}{m_{a}}\right)^{3}\left(\frac{\sigma}{10^{7}\,\text{eV}}\right)$ (31)
$\times$ $\left(\frac{1\,\mu\text{m}}{d}\right)\left(\frac{B_{t}}{1.5\times 10^{5}\,\text{Gauss}}\right)^{2}\left(\frac{\rho_{d}}{0.3\,\text{GeV}\rm cm^{-3}}\right).$
We show in Fig.(3) the excluded region of $g_{a\gamma\gamma}$ and $m_{a}$ indicated by dashed line when we do not observe $\Delta T$ equal to or larger than $50\text{mK}(t_{ob}/1\text{s})$ with observation time $t_{ob}=1$s. In the figure, we have roughly depicted the excluded region by CAPPCAPP1 ; CAPP2 ; CAPP3 . We also show the lines associated with the models of QCD axion.
Figure 3: When we do not observe $\Delta T$ larger than $5\text{mK}(t_{ob}/1\text{s})$ with observation time $t_{ob}=1$s, the value of $g_{a\gamma\gamma}$ above the dashed line denoted with $0.5\text{mK}(\tau/10\text{ms})$ is excluded. We roughly show excluded region obtained by CAPP.
Therefore, we find that a large time constant $\tau>1\,\text{s}$ is preferable to observe a measurable temperature increase $\Delta T$. To achieve such a large time constant, it is needed to minimize thermal conductance $G$ using possibly small copper lead wire and pedestal. In order to measure the temperature rise of the order of 11mK, we may adopt standard ruthenium oxide resistance thermometer, which is used for the experiment of quantum Hall effect even at low temperature $\sim 20$mK.
@@
To detect axion, we measure the increase of the temperature of material caused by the absorption of amplified but still much weak microwaves generated by the axion. So it is favorable to use material with much small heat capacity and small thermal conductance G of the material connecting to heat bath. In the paper we have proposed two dimensional electrons forming quantum Hall states as such a material. But the idea is also applicable for materials which are able to absorb the microwaves and have sufficiently small heat capacities for the temperature to increase. To find such a material is a subject in near future.
@@
Although the detailed method how we measure the temperature rise is beyond the scope of this work, the predicted increase offers a direct and sensitive means for detecting the axion and determining its mass.